OpenStudy (anonymous):

i need work shown. find the average of 9,342; 9,107 ; 10,005 ; 9,843 and 9,273?

OpenStudy (anonymous):

Hey, Average means (sum of observations)/ (number of observations) So , in here, <av> = (9342+ 9107 + 10005 + 9843 + 9273) /5 =47570/5 = 9514 Hope it makes sense :-)

OpenStudy (anonymous):

Got it???

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

My pleasure :-)